# Late-time transition of $M_{B}$ inferred via neural networks

### v1

@@ III Reconstruction Methdology
In this section we discuss the data sets used in this analysis together with our training and validation strategy for structuring the ANN system. Finally, we lay out the physical framework on which we perform this model-independent constraint analysis on the evolution of $M_{B}$.
III.1 Observational Data sets
@@
As mentioned in the introduction, in this work, we consider two sets of observational data. We consider the latest Pantheon+ compilation for SNIa observations Brout:2021mpj ; Riess:2021jrx ; Scolnic:2021amr from 1701 light curves that represent 1550 distinct SNIa spanning the redshift range $z<2.3$. This data set comprises apparent magnitude measurements $m$ with their associated statistical uncertainties, tabulated at different redshifts. It also includes a covariance matrix detailing the systematic errors or correlations in the measurements.
We also take into account the latest data set of 32 CC $H(z)$ measurements Stern:2009ep ; Moresco:2012jh ; Moresco:2016mzx ; Borghi:2021rft ; Ratsimbazafy:2017vga ; Moresco:2015cya ; Zhang:2012mp , along with the full covariance matrix that includes systematic and calibration errors, as reported in Ref. Moresco:2020fbm . These data do not assume any particular cosmological model, but depend on the differential ages technique between galaxies, covering the redshift range up to $z\sim 2$.
III.2 ANN Training and Validation
@@
Following the prescription outlined in Sec. II, we train our neural network using the Pantheon+ SNIa dataset Brout:2021mpj ; Riess:2021jrx ; Scolnic:2021amr . To optimize the network’s performance, we split the dataset into training (70%) and validation (30%) sets. To incorporate the covariance matrix of the Pantheon+ dataset into the training algorithm, the $\chi^{2}$ loss function, defined in Eq. (2) is minimized. During training, the number of iterations is set to 30,000 at every epoch where a batch size of 32 was considered. The covariance sub-matrices equivalent for every training batch were carefully chosen from the actual data ensuring that these sub-matrices are positive semi-definite.
@@
III.3 Theoretical Framework
@@
$d_{L}(z)=c(1+z)\int_{0}^{z}\frac{d\tilde{z}}{H(\tilde{z})}.$ (3)
@@
$m(z)-M_{B}=5\log_{10}\left[\frac{d_{L}(z)}{\text{1 Mpc}}\right]+25\,.$ (4)
From Eq. (4), we can rewrite the luminosity distance as,
$d_{L}(z)=10^{\frac{1}{5}\left[m(z)-M_{B}-25\right]}.$ (5)
@@
$d_{L}^{\prime}(z)=\frac{\log(10)}{5}10^{-\frac{M_{B}}{5}}m^{\prime}(z).$ (6)
Combining Eq. (5) and Eq. (6) with Eq. (3), we can express the Hubble parameter as,
$H(z)=\frac{c(1+z)^{2}}{(1+z)d_{L}^{\prime}(z)-d_{L}(z)}.$ (7)
@@
Figure 3: ANN reconstruction of the Pantheon+ apparent magnitudes $m(z)$ at the CC redshifts $(z_{\text{CC}})$.
Figure 4: Normalized covariance matrix between ANN reconstruction of the Pantheon+ SN-Ia apparent magnitudes $m(z)$ at the CC redshifts $(z_{\text{CC}})$.
IV Results and Discussions
@@ IV.1 Constraints on $M_{B}$
For these $n$-reconstructed $m$ and $m^{\prime}$, one can write down the corresponding Hubble parameter as a function of $M_{B}$ utilizing Eq. (7). Finally, to obtain the model-independent constraints on $M_{B}$ we define this $\chi^{2}$ function,
$\chi^{2}\left\{M_{B}\right\}=\sum_{i,j}\left[H_{\text{CC}}(z_{i})-H_{\text{ANN}}(z_{i})\right]^{\text{T}}\Sigma^{-1}_{ij}\left[H_{\text{CC}}(z_{j})-H_{\text{ANN}}(z_{j})\right]\,.$ (8)
@@
$\log\mathcal{L}\left\{M_{B}\right\}=-\frac{\chi^{2}}{2}-\frac{1}{2}\log|\Sigma|-\frac{N}{2}\log(2\pi)\,,$ (9)
@@
There have been similar efforts to compute $M_{B}$ from different combinations of different cosmological observations Camarena:2019rmj ; Dinda:2022jih ; Banerjee:2023evd ; Favale:2023lnp ; Mukherjee:2021epjc ; Shah:2024slr ; Gomez-Valent:2021hda ; Mukherjee:2021kcu . For a comparative analysis, we mention some of the previous results, for a given set of observational data, in Table 1. Finally, a whisker plot is shown with the available results in Fig. 6 for better illustration. We find that our results are consistent with the constraints on $M_{B}$ obtained from almost all the previous analyses, particularly when similar data sets are taken into consideration Mukherjee:2021epjc ; Dinda:2022jih ; Mukherjee:2021kcu . It deserves to mention that the constraints on $M_{B}$ are tighter when the BAO observations are combined with the CC and SNIa data. This is because the uncertainty associated with the BAO data set is significantly smaller in comparison to the CC data set. However, in this work, we refrain from including the BAO data sets because there are apprehensions regarding their dependence towards some fiducial cosmological model. So, we undertake this analysis utilizing only the SNIa and CC data sets to ensure a complete model-independent prescription.
@@
Camarena & MarraCamarena:2019rmj Cosmography $\alpha_{\text{BAO}}$ + $r_{d}^{\text{CMB}}$ + Pantheon $-19.401\pm 0.027$
@@
Mukherjee & MukherjeeMukherjee:2021kcu Gaussian Process CC + Pantheon $\mathbf{-19.387\pm 0.060}$
Mukherjee & BanerjeeMukherjee:2021epjc Gaussian Process CC + Pantheon $\mathbf{-19.360\pm 0.059}$
@@
Dinda & BanerjeeDinda:2022jih Gaussian Process CC + Pantheon $\mathbf{-19.384\pm 0.052}$
@@
Benisty et al.Benisty:2022psx Neural Networks $\alpha_{\text{BAO}}$ + $r_{d}^{\text{CMB}}$ + Pantheon $-19.38\pm 0.20$
@@
Gómez-ValentGomez-Valent:2021hda Index of Inconsistency CC + BAO + Pantheon $-19.374\pm 0.080$
@@
Favale et al.Favale:2023lnp Gaussian Process CC + Pantheon+ $-19.344^{+0.116}_{-0.090}$
@@
Banerjee et al.Banerjee:2023evd Gaussian Process CC + $r_{\text{BAO}}$ + Pantheon+ ($\Omega_{k}\neq 0$) $-19.404^{+0.099}_{-0.104}$
Shah et al.Shah:2024slr Neural Networks Pantheon + $\alpha_{\text{BAO}}$ + $r_{d}^{\text{CMB}}$ $-19.394^{+0.018}_{-0.017}$
@@
This ${}^{\star}$ Neural Networks CC & Pantheon+ $\mathbf{-19.353^{+0.073}_{-0.078}}$
@@
Figure 6: Whisker plot showing a comparison between the model-independent constraints on $M_{B}$ obtained in this work vs those present in the existing literature. The purple region corresponds to the $1-\sigma$ model independent constraint $M_{B}=-19.214\pm 0.037$, as reported in Ref. Riess:2020fzl ; Riess:2021jrx inferred from the Pantheon and SH0ES data sets.
IV.2 Data-driven transition of $M_{B}$
We now investigate the possibility of a redshift evolution in $M_{B}$ by considering two types of binning methods, applied to the adopted $H(z)$ CC data which was already used in Sec. IV.1. The first binning method is the so-called cumulative binning technique Hu:2022kes , in which the next redshift bin has one more high-redshift $H(z)$ data point with respect to the previous redshift bin. To be able to run the MCMC analysis, in our first cumulative redshift bin, we considered the first six low-redshift $H(z)$ data points up to $z_{\mathrm{max}}=0.1993$. The derived values of $M_{B}(z_{\mathrm{max}})$ from the cumulative binning method are shown in Fig. 7, where $M_{B}(z_{\mathrm{max}})$ denotes the value of $M_{B}$ inferred from the $H(z)$ data set with maximal redshift $z_{\mathrm{max}}$.
[…]
@@ IV.2 Data-driven transition of $M_{B}$
Figure 7: Predictions of the supernovae apparent magnitudes: $M_{B}(z_{\mathrm{max}})$ by adopting cumulative binning, where $M_{B}(z_{\mathrm{max}})$ is the derived value of $M_{B}$ by considering CC $H(z)$ data up to $z_{\mathrm{max}}$ (left panel), and $M_{B}(\bar{z})$ by adopting the redshift layer binning, where $M_{B}(\bar{z})$ is the derived value of $M_{B}$ by considering CC $H(z)$ data within that redshift layer with a mean redshift $\bar{z}$ (right panel). The purple region corresponds to the $1-\sigma$ model independent constraint $M_{B}=-19.214\pm 0.037$, as reported in Ref. Riess:2020fzl ; Riess:2021jrx inferred from the Pantheon and SH0ES data sets.
V Conclusion
@@
In the usual practice, when one estimates cosmic distances as a function of the scale factor and its derivatives, the luminosity of type Ia supernovae is considered a standard candle, which is of utmost importance to understanding the accelerating expansion of the Universe. In this work, we have tried to investigate if this assumption is as educated as it sounds, by constraining the peak absolute magnitude, $M_{B}$, of SNIa, in a cosmological model-independent way. For this exercise, we have utilized two distinct sets of observational data, namely the Pantheon+ SNIa compilation and the cosmic chronometer Hubble parameter measurement, as described in Sec. III.1
@@
In this work we explore the nature of the absolute magnitude parameter through the use of ANNs with a focus on obtaining a model-independent value of this parameter, as well as an indication of whether the parameter expresses a statistically relevant evolution. Using Eq. 7, we show how the Pantheon+ sample can be used to determine the Hubble parameter at different redshift points. This is then used in conjunction with the CC Hubble parameter observations to constrain the value of $M_{B}$ in a model-independent way using the log-likelihood defined in Eq. 9. As shown in Table 1, the final constraint on $M_{B}$ is competitive with other literature values, and moreover, this approach had the advantage that it does not depend on a fiducial cosmological model.
@@
Since the focus is on constraining $M_{B}$ in a model-agnostic way, we do not consider any parametric or theoretical form between $d_{L}$ and $H(z)$ at the outset. Hence, once the observed apparent magnitudes $m$ of the SNIa are given, we proceed to reconstruct $m(z)$ a function of the redshift employing Artificial Neural Networks (Sec. III.2). Utilizing the equations in Sec. III.3, the analogous Hubble parameter, inferred from SN-Ia via Eq. (7), can in principle be written as a function of its peak absolute magnitude $M_{B}$. Finally, having defined the $\chi^{2}$ function for $M_{B}$ (Eq. (8)), we obtain constraints by minimizing the negative log-likelihood in Eq. (9). The result we get is
$M_{B}=-19.353^{+0.073}_{-0.078}\,.$ (10)
@@
Furthermore, we proceed to test the evolution of the peak absolute magnitude $M_{B}$ as a function of redshift. For this exercise, two types of binning methods have been adopted while working with the CC data. In the first one, we used the cumulative method, in which a redshift bin has one more data point with respect to the previous bin and the results are shown in Fig. 7. In the second one, we considered four distinct binning of redshifts, where each one has a mean redshift $\bar{z}$, and the results are presented in Table 2. In both cases, one can notice a transition in the value of $M_{B}$ at $z\approx 1$.
@@
More broadly, previous works in the literature open the possibility of a transition redshift over which some parameters may suffer a change in value which includes the $M_{B}$ parameter. Ref. Akarsu:2023mfb ; Paraskevas:2024ytz points to a sign switch from negative to positive in the cosmological constant, while Ref. Perivolaropoulos:2021bds indicates a value transition in $M_{B}$ albeit at a much closer distance. Indeed, even a regular MCMC albeit with a binning procedure will produce a variety of $H_{0}$ values as shown in Refs. Colgain:2023bge ; Colgain:2022tql which may ultimately be interpreted as a variation in the $M_{B}$ parameter.
@@
Indications in the literature have pointed to a variation in the value of $M_{B}$ Efstathiou:2021ocp ; Camarena:2022iae ; Camarena:2021jlr which may represent the underlying physical nature of the tension in terms of astrophysics. The problem of the source and characterization of the $H_{0}$ tension may express other important features relevant for further study. In future work, we hope to apply this new approach to assessing the evolution profile of other important parameters that contribute to constraints on the value of the Hubble constant.
Acknowledgements.

### v2

@@ 3 Reconstruction Methodology
In this section, we discuss the data sets used in this analysis together with our training and validation strategy for structuring the ANN system. Finally, we lay out the physical framework on which we perform this model-independent constraint analysis on the evolution of $M_{B}$.
3.1 Observational Data sets
@@
As mentioned in the introduction, in this work, we consider two sets of observational data. We consider the latest Pantheon+ compilation for SNIa observations [128, 129, 130] from 1701 light curves that represent 1550 distinct SNIa spanning the redshift range $z<2.3$. This data set comprises apparent magnitude measurements $m$ with their associated statistical uncertainties, tabulated at different redshifts. It also includes a covariance matrix detailing the systematic errors or correlations in the measurements. Recent investigations with Pantheon+ have highlighted concerns, including a possible 5% overestimation of uncertainties in the covariance matrix [131] and the potential presence of an unaddressed systematic effect—volumetric redshift scatter—that could suggest a physical shift in the absolute magnitude of SNIa at redshift $z=0.005$ [132].
We also take into account the latest data set of 32 CC $H(z)$ measurements [133, 134, 135, 136, 137, 138, 139], along with the full covariance matrix that includes systematic and calibration errors, as reported in Ref. [140]. These data do not assume any particular cosmological model, but depend on the differential ages technique between galaxies, covering the redshift range up to $z\sim 2$. The CC method, although a new emerging cosmological probe that provides model-independent estimates of the Universe’s expansion history, its reliability is often debated within the cosmology community [14]. Major large-scale cosmological surveys like the SDSS, DESI, or DES, prioritize other observational techniques like SNIa, BAO, and galaxy clustering over the CC method. Its reliance on the age of the Universe, being an integrated quantity of $H(z)$, faces limitations in statistical power and susceptibility to systematic errors, such as the calibration of Stellar Population Synthesis (SPS) models and the accurate selection of CC samples to minimize contamination [141]. Although we incorporate this data into our analysis, its precision and accuracy are compromised by uncertainties in metallicity, biases from star formation history, and systematic errors from spectral models [142]. These factors introduce potential biases and interpretational challenges (see Ref. [88, 143, 67] for the possible impact of CC systematics on cosmological results), reducing its robustness and ability to independently drive our conclusions. Therefore, our findings are based on the assumption of this data’s reliability, recognizing the need for further validation and scrutiny.
3.2 ANN Training and Validation
@@
Following the prescription outlined in Sec. 2, we train our neural network using the Pantheon+ SNIa dataset [128, 129, 130]. To optimize the network’s performance, we split the dataset into training (70%) and validation (30%) sets. To incorporate the covariance matrix of the Pantheon+ dataset into the training algorithm, the $\chi^{2}$ loss function, defined in Eq. (2.2) is minimized. During training, the number of iterations is set to 30,000 at every epoch where a batch size of 32 was considered. The covariance sub-matrices equivalent for every training batch were carefully chosen from the actual data ensuring that these sub-matrices are positive semi-definite.
@@
Figure 3: ANN reconstruction of the Pantheon+ apparent magnitudes $m(z)$ (left panel) and its corresponding derivatives $m^{\prime}(z)$ [right panel] at the CC redshifts $(z_{\text{CC}})$.
Figure 4: Normalized covariance matrices between ANN reconstruction of the Pantheon+ SN-Ia apparent magnitudes $m(z)$ and its corresponding derivatives $m^{\prime}(z)$ at the CC redshifts $(z_{\text{CC}})$.
3.3 Theoretical Framework
@@
$d_{L}(z)=c(1+z)\int_{0}^{z}\frac{d\tilde{z}}{H(\tilde{z})}.$ (3.1)
@@
$m(z)-M_{B}=5\log_{10}\left[\frac{d_{L}(z)}{\text{1 Mpc}}\right]+25\,.$ (3.2)
From Eq. (3.2), we can rewrite the luminosity distance as,
$d_{L}(z)=10^{\frac{1}{5}\left[m(z)-M_{B}-25\right]}.$ (3.3)
@@
$d_{L}^{\prime}(z)=\frac{\log(10)}{5}10^{-\frac{M_{B}}{5}}m^{\prime}(z).$ (3.4)
Combining Eq. (3.3) and Eq. (3.4) with Eq. (3.1), we can express the Hubble parameter as,
$H(z)=\frac{c(1+z)^{2}}{(1+z)d_{L}^{\prime}(z)-d_{L}(z)}.$ (3.5)
@@
4 Results and Discussions
@@ 4.1 Constraints on $M_{B}$
For these $n$-reconstructed $m$ and $m^{\prime}$, one can write down the corresponding Hubble parameter as a function of $M_{B}$ utilizing Eq. (3.5). Finally, to obtain the model-independent constraints on $M_{B}$ we define this $\chi^{2}$ function,
$\chi^{2}\left\{M_{B}\right\}=\sum_{i,j}\left[H_{\text{CC}}(z_{i})-H_{\text{ANN}}(z_{i})\right]^{\text{T}}\Sigma^{-1}_{ij}\left[H_{\text{CC}}(z_{j})-H_{\text{ANN}}(z_{j})\right]\,.$ (4.1)
@@
$\log\mathcal{L}\left\{M_{B}\right\}=-\frac{\chi^{2}}{2}-\frac{1}{2}\log|\Sigma|-\frac{N}{2}\log(2\pi)\,,$ (4.2)
@@
Camarena & Marra[144] Cosmography $\alpha_{\text{BAO}}$ + $r_{d}^{\text{CMB}}$ + Pantheon $-19.401\pm 0.027$
@@
Mukherjee & Mukherjee[118] Gaussian Process CC + Pantheon $\mathbf{-19.387\pm 0.060}$
Mukherjee & Banerjee[61] Gaussian Process CC + Pantheon $\mathbf{-19.360\pm 0.059}$
@@
Dinda & Banerjee[115] Gaussian Process CC + Pantheon $\mathbf{-19.384\pm 0.052}$
@@
Benisty et al.[68] Neural Networks $\alpha_{\text{BAO}}$ + $r_{d}^{\text{CMB}}$ + Pantheon $-19.38\pm 0.20$
@@
Gómez-Valent[145] Index of Inconsistency CC + BAO + Pantheon $-19.374\pm 0.080$
@@
Favale et al.[122] Gaussian Process CC + Pantheon+ $-19.344^{+0.116}_{-0.090}$
@@
Banerjee et al.[119] Gaussian Process CC + $r_{\text{BAO}}$ + Pantheon+ ($\Omega_{k}\neq 0$) $-19.404^{+0.099}_{-0.104}$
Shah et al.[111] Neural Networks Pantheon + $\alpha_{\text{BAO}}$ + $r_{d}^{\text{CMB}}$ $-19.394^{+0.018}_{-0.017}$
@@
This work Neural Networks CC & Pantheon+ $\mathbf{-19.353^{+0.073}_{-0.078}}$
@@
Figure 6: Whisker plot showing a comparison between the model-independent constraints on $M_{B}$ obtained in this work vs those present in the existing literature. The purple region corresponds to the $1-\sigma$ model independent constraint $M_{B}=-19.214\pm 0.037$, as reported in Ref. [146, 129] inferred from the Pantheon and SH0ES data sets.
There have been similar efforts to compute $M_{B}$ from different combinations of cosmological observations [144, 115, 119, 122, 61, 111, 145, 118]. For a comparative analysis, we mention some of the previous results, for a given set of observational data, in Table 1. Finally, a whisker plot is shown with the available results in Fig. 6 for better illustration. We find that our results are consistent with the constraints on $M_{B}$ obtained from almost all the previous analyses, particularly when similar data sets are taken into consideration [61, 115, 118]. It deserves to be mentioned that the constraints on $M_{B}$ are tighter when the BAO observations are combined with the CC and SNIa data. This is because the uncertainty associated with the BAO data set is significantly smaller in comparison to the CC data set. However, in this work, we refrain from including the BAO data sets because there are apprehensions regarding their dependence towards some fiducial cosmological model. So, we undertake this analysis utilizing only the SNIa and CC data sets to ensure a complete model-independent prescription.
@@
Figure 7: Predictions of the supernovae absolute magnitudes: $M_{B}(z_{\mathrm{max}})$ by adopting cumulative binning, where $M_{B}(z_{\mathrm{max}})$ is the derived value of $M_{B}$ by considering CC $H(z)$ data up to $z_{\mathrm{max}}$ (left panel), and $M_{B}(\bar{z})$ by adopting the redshift layer binning, where $M_{B}(\bar{z})$ is the derived value of $M_{B}$ by considering CC $H(z)$ data within that redshift layer with a mean redshift $\bar{z}$ (right panel). The purple region corresponds to the $1-\sigma$ model independent constraint $M_{B}=-19.214\pm 0.037$, as reported in Ref. [146, 129] inferred from the Pantheon and SH0ES data sets.
4.2 Data-driven transition of $M_{B}$
We now investigate the possibility of a redshift evolution in $M_{B}$ by considering two types of binning methods, applied to the adopted $H(z)$ CC data which was already used in Sec. 4.1. The first binning method is the so-called cumulative binning technique [147, 148], in which the next redshift bin has one more high-redshift $H(z)$ data point with respect to the previous redshift bin. To be able to run the MCMC analysis, in our first cumulative redshift bin, we considered the first six low-redshift $H(z)$ data points up to $z_{\mathrm{max}}=0.1993$. The derived values of $M_{B}(z_{\mathrm{max}})$ from the cumulative binning method are shown in Fig. 7, where $M_{B}(z_{\mathrm{max}})$ denotes the value of $M_{B}$ inferred from the $H(z)$ data set with maximal redshift $z_{\mathrm{max}}$.
[…]
5 Conclusion
@@
In the standard practice, when estimating the cosmic distances as a function of the scale factor and its derivatives, the intrinsic luminosity of Type Ia supernovae is considered a standard candle. This assumption is of utmost importance in understanding the accelerated expansion of the Universe. In this study, we have tried to investigate if this assumption is as educated as it sounds, by constraining the peak absolute magnitude, $M_{B}$, of SNIa, in a cosmological model-independent way. For this exercise, we have utilized two distinct sets of observational data, namely the Pantheon+ SNIa compilation and the cosmic chronometer Hubble parameter measurement, as described in Sec. 3.1.
@@
In this work, we have explored the nature of the absolute magnitude parameter through the use of ANNs with a focus on obtaining a model-independent value of this parameter, as well as assessing whether this parameter exhibits a statistically relevant evolution. We demonstrate how the Pantheon+ sample can be utilized, through Eq. 3.5, to determine the Hubble parameter at different redshift points. This information is then used in conjunction with the CC Hubble parameter observations to constrain the value of $M_{B}$ in a model-independent manner, employing the log-likelihood defined in Eq. 4.2. Our results, summarized in Table 1, indicate that the final constraint on $M_{B}$ is competitive with values in the existing literature. Additionally, this approach had the advantage of not relying on a fiducial cosmological model.
@@
Since the focus is on constraining $M_{B}$ in a model-agnostic way, we do not consider any parametric or theoretical form between $d_{L}$ and $H(z)$ at the outset. Instead, given the observed apparent magnitudes $m$ of the SNIa, we proceed to reconstruct $m(z)$ a function of the redshift employing Artificial Neural Networks (Sec. 3.2). Utilizing the equations in Sec. 3.3, the analogous Hubble parameter, inferred from SN-Ia via Eq. (3.5), can in principle be written as a function of its peak absolute magnitude $M_{B}$. Finally, having defined the $\chi^{2}$ function for $M_{B}$ (Eq. (4.1)), we obtain constraints by minimizing the negative log-likelihood in Eq. (4.2). The result we get is
$M_{B}=-19.353^{+0.073}_{-0.078}\,.$ (5.1)
@@
Furthermore, we proceed to test the evolution of the peak absolute magnitude $M_{B}$ as a function of redshift. For this exercise, two types of binning methods have been adopted while working with the CC data. In the first one, we used the cumulative method, in which a redshift bin has one more data point with respect to the previous bin and the results are shown in Fig. 7. In the second one, we considered four distinct binning of redshifts, where each one has a mean redshift $\bar{z}$, and the results are presented in Table 2. In both cases, one can notice a transition in the value of $M_{B}$ at $z\approx 1$. Given that both methods result in consistent transitional redshifts, this indicates that the result is likely to be a genuine result of the underlying pipeline method rather than a statistical irregularity. To further validate our reconstruction approach and analysis methodology, we have added an appendix A dedicated to a thorough test of our statistical pipeline. This includes further mock simulations specifically crafted to assess the estimation of $M_{B}$ and to detect potential redshift variations in $M_{B}$ as a function of $z$. We considered two different scenarios for injecting $M_{B}$ to create the Pantheon+-like mock apparent magnitude $m(z)$ data. This extensive testing enhances the confidence in the robustness and reliability of our method.
@@
More broadly, previous works in the literature suggest the possibility of a transition redshift at which certain parameters, including the $M_{B}$ parameter, may suffer a change in value. Ref. [150, 151] points to a sign switch in the cosmological constant from positive to negative, while Ref. [152, 153] indicates a transition or evolution in the value of $M_{B}$ albeit at a much closer distance. Indeed, even a regular MCMC albeit with a binning procedure will produce a variety of $H_{0}$ values as shown in Refs. [154, 25, 155, 156, 157, 158] some of which may ultimately be interpreted as a variation in the $M_{B}$ parameter.
@@
There are indications towards a variation in the value of $M_{B}$ present in the existing literature [114, 113, 112], which may represent the underlying physical nature of the tension in terms of astrophysics. Understanding the source and characterizing the $H_{0}$ tension may express other important features relevant for further study. In future work, we hope to apply novel machine learning approaches to assessing the evolution profile of other important, rather critical, parameters that significantly contribute to or influence the constraints on the value of the Hubble constant.
Appendix A Method Validation with Mock Simulation
Figure 8: Reconstruction of the simulated mock Pantheon+ like SNIa apparent magnitudes $m(z)$: for Case (i) a constant $M_{B}=-19.4$ for all redshifts $z$ [in left panel] and Case (ii) with a transition viz. $M_{B}=-19.4$ for $z\geq 0.5$ and $M_{B}=-19.2$ for $z<0.5$. [in right panel].
To validate our pipeline, we use a fiducial $\Lambda$CDM cosmology with parameters $\Omega_{m0}=0.3$ and $H_{0}=70\,\text{km}\,\text{Mpc}^{-1}\text{s}^{-1}$. We generate mock data using the Pantheon+ compilation and CC Hubble parameter measurements. For the mock Pantheon+ data, we perform two types of injections for the absolute magnitude value: (i) a constant $M_{B}=-19.4$ for all redshifts $z$, and (ii) a transition where $M_{B}=-19.4$ for $z\geq 0.5$ and $M_{B}=-19.2$ for $z<0.5$. To account for uncertainties, we introduce random Gaussian noise corresponding to 2% of the magnitude values for both mock $m(z)$ and $H(z)$ respectively.
[…]
Figure 9: Predictions for the mock SNIa absolute magnitudes: $M_{B}(z_{\mathrm{max}})$ by adopting cumulative binning (left panel), and $M_{B}(\bar{z})$ by adopting the redshift layer binning (right panel). The top panel shows Case (i) injection of a constant $M_{B}=-19.4$ for all $z$, whereas the bottom panel shows Case (ii) injection with a transition viz. $M_{B}=-19.4$ for $z\geq 0.5$ and $M_{B}=-19.2$ for $z<0.5$. The blue dashed line shows the injected $M_{B}$ values. The purple region corresponds to $M_{B}=-19.214\pm 0.037$ [146, 129].
Second, we aim to investigate the potential redshift evolution of $M_{B}$ through a null hypothesis test, utilizing both cumulative and layered binning methods as outlined in Sec. 4.2. In Fig. 9, we present the predictions of simulated SNIa absolute magnitudes for both Case (i) and Case (ii) mock $M_{B}$ injections. The left panel displays the predictions of $M_{B}(z_{\mathrm{max}})$ using cumulative binning, where $M_{B}(z_{\mathrm{max}})$ is the value derived from mock $H(z)$ data up to $z_{\mathrm{max}}$. The right panel shows the constraints on $M_{B}(\bar{z})$ obtained through redshift layer binning, where $M_{B}(\bar{z})$ is derived from mock $H(z)$ data within a redshift layer centred around a mean redshift $\bar{z}$.
[…]
Acknowledgments